# Investigation of Time-Frequency Feature Combinations with Histogram Layer Time Delay Neural Networks

### v1

@@ Abstract
deep learning, underwater acoustic classification, histogram layers, feature engineering
@@ I Introduction
Images can be decomposed into statistical and structural texture features [11, 12, 13]. Spectrograms can also be represented through these two texture types. Statistical texture features provide information about the acoustic texture [14], using statistical measures that usually capture the distribution of textures. Structural texture features, on the other hand, are the spatial patterns of textures [15]. Models like CNNs and time delay neural networks (TDNNs), are very effective at capturing spatial and temporal arrangements, or structural textures in imagery such as spectrograms [13]. However, these models may not fully take into account the statistical features within spectrograms.
Figure 1: Overall workflow of this work. The audio segments are first partitioned into individual segments. Each audio segment is then used to compute an individual feature of varying time and frequency resolutions. We introduce an adaptive padding layer to make each spectrogram the same size. The features are then concatenated along the channel dimension and processed by the HLTDNN model for classification.
@@
The representation of audio signals is a difficult problem yet crucial to the success of machine learning algorithms [17]. Related work [18] has demonstrated that combining three types of spectrograms, mel-frequency spectrograms, gammatone-frequency spectrograms (Cochleagrams), and continuous wavelet transforms (CWT), can improve the detection of speech deficits in cochlear implant users. The task of acoustic feature engineering, particularly for underwater environments, still remains challenging [19]. While a prior study showed feature fusion to be effective in improving UATR performance [20], it focused on a smaller UATR dataset, ShipsEar [21]. A related effort [22] used a larger dataset, DeepShip [23], to show that feature fusion could improve performance, but did not investigate the impact of feature selection for three spectrogram features used. Additionally, spectrograms for the inputs are restricted to be the same size and cannnot use different hyperparameters (e.g., window length, hop length) for each spectrogram which may unintentionally restrict the unique information captured by each feature.
@@ II-A Review of the HLTDNN architecture and the histogram layer operation
II-B Feature combination implementation
We propose merging different spectrogram features to potentially improve the representational power of the input data passed into the model. Given a total of $\mathcal{M}$ unique time-frequency features, the total number of unique combinations, $\mathcal{N}$, can be computed as $\mathcal{N}=2^{\mathcal{M}}-1$. For our work, six different time-frequency features are combined to generate 63 distinct combinations ($12^{6}-$ 1). To prepare these combination of features for input into the model, an adaptive zero padding layer is incorporated into the spectrograms. During adaptive padding, since features do not share the same height and width dimensions, they are resized to a common dimension, as to be combined without losing their time-frequency resolution by calculating the required padding for each side of a feature map to match the largest height and width among all features. Padding is applied symmetrically to maintain the feature’s original structure and position within its new dimensions. Once all features are resized to a uniform dimension, they are combined by concatenating the feature maps along the channel dimension. The overall workflow is shown in Fig. 1.
III Experimental Setup
In this work, the DeepShip sonar dataset is used [23]. This dataset has a collection of real-world sonar recordings with 265 different ships across four classes (tanker, tug, passengership, and cargo), compiled from 47 hours and 4 minutes of recordings. Recordings are captured under various sea states and noise levels. Audio signals are resampled to a frequency of 16 kHz and then segmented into intervals of three seconds. Next, six time-frequency based features are extracted: mel-frequency spectrogram (MS), mel-frequency cepstral coefficients (MFCCs), short-time fourier transform (STFT), gammatone-frequency cepstral coefficients (GFCC), constant-q transform (CQT), and variable-q transform (VQT). Studies have demonstrated the effectiveness of these features in classifying underwater acoustic signals [23, 25, 16]. For the extraction process, the window and hop length are set at 250 ms and 64 ms, respectively. Specifically, MS uses 44 Mel filter banks. In the case of MFCCs, there are 16 mel-frequency cepstral coefficients. Meanwhile, STFT has 48 frequency bins. For the GFCC, CQT, and VQT features, there are 64 frequency bins.
The dataset was partitioned into training, validation, and test sets with respective ratios of 70%, 15%, and 15% based on [16]. A total of 56,468 segments were generated, distributed as 38,523 for training, 9,065 for validation, and 8,880 for testing. To prevent data leakage, segments derived from a single recording were exclusively allocated to the same partition. This ensures if a recording is chosen for training, all its segments are also chosen solely for training purposes. The HLTDNN model was evaluated with each combination of features across three runs of random initialization. The configuration for the experimental parameters was as follows: Adagrad optimizer with a learning rate of 0.001 and training batch size of 128. A dropout rate of 0.5 was applied preceding the output classification layer to mitigate overfitting. Additionally, early stopping was implemented to stop training if no improvement in validation loss is observed after 15 epochs. The maximum epoch number was set to 150. Following [16], the number of bins was set to 16 for the HLTDNN model. Experiments were completed on an A100 GPU.
IV Results and Discussion
@@
(a) Best: $66.17\pm 1.10$%
(b) MFCC: $59.34\pm 3.16$%
Figure 2: Comparison of confusion matrix results for the best combination of features and MFCC alone.
(a) Best (FDR: $61.26\pm 4.52$)
(b) MFCC (FDR: $29.04\pm 2.49$)
Figure 3: t-SNE results for the best feature combination and MFCC. Colors show the four classes of ships. The average log Fisher Discriminant Ratio (FDR) (±$\pm 1\sigma$) is also shown. Higher scores show more compact and better separated classes. The best random seed is used for each feature.
@@
(a) Original VQT
(b) Original STFT
(c) Original GFCC
(d) Best CAM
(e) Best CAM
(f) Best CAM
(g) Original MFCC
(h) Best CAM
(i) MFCC CAM
Figure 4: Class Activation Maps (CAM) using the best feature combination and MFCC alone for the Cargo class. The third row compares the CAM overlay on the MFCC for both of the models.

### v2

# College Station, TX, USA

@@ Abstract
deep learning, underwater acoustic, histogram layers, feature engineering
@@ I Introduction
Images can be decomposed into statistical and structural texture features [11, 12, 13]. Spectrograms can also be represented through these two texture types. Statistical texture features provide information about the texture [14], using statistical measures that usually capture the distribution. Structural texture features, on the other hand, are the spatial patterns of textures [15]. Models like and time delay neural networks (TDNNs) are very effective at capturing spatial and temporal arrangements, or structural textures in imagery such as spectrograms [13]. However, these models may not fully take into account the statistical features within spectrograms.
Figure 1: Overall workflow. The audio segments are first partitioned into individual segments. Each audio segment is then used to compute an individual feature of varying time and frequency resolutions. We introduce an adaptive padding layer to make each spectrogram the same size. The features are then concatenated along the channel dimension and processed by the HLTDNN model for classification.
@@
The representation of audio signals is a difficult problem yet crucial to the success of machine learning algorithms [17]. Related work [18] has demonstrated that combining three types of spectrograms, mel-frequency spectrograms, gammatone-frequency spectrograms (Cochleagrams), and continuous wavelet transforms (CWT), can improve the detection of speech deficits in cochlear implant users. The task of acoustic feature engineering, particularly for underwater environments, still remains challenging [19]. While a prior study showed feature fusion to be effective in improving UATR performance [20], it focused on a smaller UATR dataset, ShipsEar [21]. A related effort [22] used a larger dataset, DeepShip [23], to show that feature fusion could improve performance, but did not investigate the impact of feature selection for three spectrogram features used. Additionally, spectrograms for the inputs are restricted to be the same size and cannot use different hyperparameters (e.g., window length, hop length) for each spectrogram which may unintentionally restrict the unique information captured by each feature.
@@ II-A Review of the HLTDNN architecture and the histogram layer operation
II-B Feature combination implementation
We propose merging different spectrogram features to potentially improve the representational power of the input data passed into the model. Given a total of $\mathcal{M}$ unique time-frequency features, the total number of unique combinations, $\mathcal{N}$, can be computed as $\mathcal{N}=2^{\mathcal{M}}-1$. For our work, six different time-frequency features are combined to generate 63 distinct combinations ($12^{6}-$ 1). To prepare these combinations of features for input into the model, an adaptive zero padding layer is incorporated into the spectrograms. During adaptive padding, since features do not share the same height and width dimensions, they are resized to a common dimension, as to be combined without losing their time-frequency resolution by calculating the required padding for each side of a feature map to match the largest height and width among all features. Padding is applied symmetrically to maintain the feature’s original structure and position within its new dimensions. Once all features are resized to a uniform dimension, they are combined by concatenating the feature maps along the channel dimension. The overall workflow is shown in Fig. 1.
III Experimental Setup
In this work, the DeepShip sonar dataset is used [23]. This dataset has a collection of real-world sonar recordings with 265 different ships across four classes (tanker, tug, passengership, and cargo), compiled from 47 hours and 4 minutes of recordings. Recordings are captured under various sea states and noise levels. Audio signals are uniformly sampled to a frequency of 16 kHz and then segmented into intervals of three seconds. Next, six time-frequency based features are extracted: mel-frequency spectrogram (MS), mel-frequency Cepstral coefficients (MFCCs), short-time Fourier transform (STFT), gammatone-frequency Cepstral coefficients (GFCC), constant-Q transform (CQT), and variable-Q transform (VQT). Studies have demonstrated the effectiveness of these features in classifying underwater acoustic signals [23, 25, 16]. For the extraction process, the window and hop length are set at 250 ms and 64 ms, respectively. Specifically, MS uses 44 Mel filter banks. In the case of MFCCs, there are 16 mel-frequency cepstral coefficients. Meanwhile, STFT has 48 frequency bins. For the GFCC, CQT, and VQT features, there are 64 frequency bins.
The dataset was partitioned into training, validation, and test sets with respective ratios of 70%, 15%, and 15% based on [16]. A total of 56,468 segments were generated, distributed as 38,523 for training, 9,065 for validation, and 8,880 for testing. To prevent data leakage, segments derived from a single recording were exclusively allocated to the same partition. This ensures if a recording is chosen for training, all its segments are also chosen solely for training purposes. The HLTDNN model was evaluated with each combination of features across three runs of random initialization. The configuration for the experimental parameters was as follows: Adagrad optimizer with a learning rate of 1e-3 and a batch size of 128. A dropout rate of 0.5 was applied preceding the output classification layer to mitigate overfitting. Additionally, early stopping was implemented to stop training if no improvement in validation loss is observed after 15 epochs. The maximum epoch number was set to 150. Following [16], the number of bins was set to 16 for the HLTDNN model.
IV Results and Discussion
@@
Figure 2: Comparison of confusion matrix results for (a) best combination of features (66.17 $\pm$ 1.10%) and (b) MFCC alone (59.34 $\pm$ 3.16%).
Figure 3: t-SNE results for (a) best feature combination (FDR: 61.26 $\pm$ 4.52) and (b) MFCC (FDR: 29.04 $\pm$ 2.49). Colors represent four classes of ships. The average log Fisher Discriminant Ratio (FDR) (±$\pm 1\sigma$) is also shown, with higher scores indicating more compact and better-separated classes. The best random seed is used for each feature.
@@
Figure 4: Class Activation Maps (CAM) using the best feature combination and MFCC alone for the Cargo class. The third row compares the CAM overlay on the MFCC for both models. (a) Original VQT, (b) Original STFT, (c) Original GFCC, (d) Best CAM for VQT, (e) Best CAM for STFT, (f) Best CAM for GFCC, (g) Original MFCC, (h) Best CAM, (i) MFCC CAM.